CNNBC recently reported that the mean annual cost of auto insurance is 964 dollars. Assume the standard deviation is 203 dollars. You take a simple random sample of 63 auto insurance policies. (Do not use tables unless directed to do so.) Find the probability that a single randomly selected value exceeds 969 dollars. P(X > 969) = Find the probability that a sample of size n = 63 is randomly selected with a mean that exceeds 969 dollars. P( ¯ x > 969) = Enter your answers as numbers accurate to 4 decimal places.
Answer #1

Solution :

Given that ,

mean = =$ 964

standard deviation = =$ 203

P(x >$969 ) = 1 - P(x<969 )

= 1 - P[(x -) / < (969 -964 ) /203 ]

= 1 - P(z <0.02 )

Using z table

= 1 - 0.5080

= 0.4920

probability= 0.4920

(B)

n = 63

= 964

= / n = 203 / 63 = 25.5756

P( >$969 ) = 1 - P( < $969)

= 1 - P[( - ) / < (969 -964) / 25.5756 ]

= 1 - P(z <0.20 )

Using z table

= 1 - 0.5793

= 0.4207

probability= 0.4207